Find the critical value of t for a 98​% confidence interval with df equals = 19.

1. t = [ ]? Round to two decimals as needed
Answer #1

Solution :

Given that,

Degrees of freedom = df = 19

At 98% confidence level

\alpha = 1 - 98%

\alpha =1 - 0.98 =0.02

\alpha/2 = 0.01

t\alpha/2,df = t0.01, 19 = 2.54

critical value = 2.54